OpenStudy (anonymous):

x, 3x-1, 4x+2, ... In the sequence above, each term after the first is 7 greater than the preceding term. What is the value of the either term in the sequence? The answer is 53 but why :C?

OpenStudy (anonymous):

x=x, x=(3x-1)+7

OpenStudy (anonymous):

huh... i had the formula set up weird :l but YAY thank you!

OpenStudy (anonymous):

let me dbl check tho

OpenStudy (anonymous):

i made a mistake

OpenStudy (anonymous):

x+7=3x-1 x=4 so plug it to the 3rd I think

OpenStudy (anonymous):

4, 11, 18, ... yeah that does work! thanks :D

OpenStudy (anonymous):

but why 53?

OpenStudy (anonymous):

it works that they increase by 7 but where is the 53 coming from?

OpenStudy (anonymous):

thats just the correct answer. thats the 8th # in the sequence

OpenStudy (anonymous):

i dont understand lol I just now I meet the requirement they go up by 7 :)

OpenStudy (anonymous):

no i figured it out now thanks you you. because of the formula u gave me, 4+7(8-1)=53 :3 thank you!!

OpenStudy (anonymous):

4+7(8-1)=53 where you get that formula lol ok i wont argue that I am right just as long as you are satisfied with your answer :)

OpenStudy (anonymous):

its a continuation of your formula D: because you found out that x=4

OpenStudy (anonymous):

Well what is the formula for the 8th term in terms of x?

OpenStudy (anonymous):

exactly as i said, basically. t^8=4+(8-1) t^8=53